40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Nov 25 02:04:19 2014, Jenne, Update, LSC, PRFPMI tonight - need some PRCL and MICH tuning at high arm powers HighPower_7pt3sec_Powers.pngHighPower_7pt3sec_ErrAndCtrls.pngHighPower_7pt3sec_AuxErrs.pngCARMDARMonALScommdiff_StayingAbove5_upTo175_24Nov2014.png
    Reply  Wed Nov 26 05:15:48 2014, Jenne, Update, LSC, PRFPMI tonight PRMI 100Hz osc? PRFPMI_currentErrSigs_25Nov2014.pdf
Message ID: 10736     Entry time: Wed Nov 26 05:15:48 2014     In reply to: 10734
Author: Jenne 
Type: Update 
Category: LSC 
Subject: PRFPMI tonight PRMI 100Hz osc? 

[Jenne, EricQ]

Just to get our day started right, we tweaked up the alignment of the Ygreen to the Yarm (after IR alignment), and also touched up the X beatnote alignment on the PSL table.  Ran the LSC offsets script, and then started locking. 

All of the locking tonight has been based on CARM and DARM held on ALS comm/diff, and PRMI held on REFL165.  Today, CARM was actuated using MC2.  No special reason for the switch from ETMs. The AS port is noticeably darker when using REFL165 instead of REFL33. 

Around 12:33am(ish), we were able to hold the arms at powers of about 100, for almost a minute.  The fluctuations were at least 50% of that value, but the average was pretty high.  Exciting.

Q and I tried a few times to engage the AO path while the arms were held at these high powers.  Q hopefully remembers what the gain and sign values were where we lost lock.  We didn't pursue this very far, since I was seeing the 50Hz oscillation that Koji and I saw the other day.  I increased the CARM gain from 6 to 10, and that seemed to help significantly.  Also, messing with the PRMI loops a bit helped.  Q increased the pole frequency in FM 5 for both MICH and PRCL from 2k to 3k.  While he had Foton open, he made sure that all of the LSC DoF filters use the z:p notation. 

I then did a few trials of trying to transition CARM over to normalized REFL11I.  Now that I'm typing, it occurs to me that I should have checked REFL11's demod phase.  Ooops.  Anyhow, using the phase that was in there, I turned on a cal line pushing on ETMs CARM, and found that using -0.002*REFL11I / (TRX + TRY) was the right set of elements.  I also put an offset of 0.05 into the CARM CESAR RF place, and started moving.  I tried several times, but never got past about 30% normalized REFL11 and 70% ALS comm. 

During these trials, Q and I worked also on tweaking up the PRMI lock.  As mentioned last night, PRCL FM3 eats too much phase (~30deg at 100Hz!), so I don't turn that on ever.  But, I do turn on FM1 (which is new tonight), FM2, 6, 8 and 9.  FM8 is a flat gain of 0.6 that I use so I can have higher gain to make acquisition faster, but immediately turn the gain down to keep the loop in the center of the phase bubble.  MICH needed a lowpass, so in addition to FM2, I am now also triggering FM 8, which is a 400Hz lowpass that was already in there. 

Now, my MICH gain is 2.4, with +0.75*REFL165Q, and PRCL gain is -0.02 with -3*REFL165I.  Triggering for both MICH and PRCL is 1*POP22I + 5*TRX with 50 up, 0.1 down. 

In my latest set of locks, I have been losing lock semi-regularly due to a 100Hz oscillation in either the PRCL or MICH loops.  If I watch the spectra, most times I take a step in CARM offset reduction, I get a broad peak in both the MICH and PRCL error signals.  Most of the time, I stay locked, and the oscillation dies away.  Sometimes though it is large enough to put me out of lock.  I'm not sure yet where this is coming from, but I think it's the next thing that needs fixing.

Here is a shot of the spectra just as one of these 100Hz oscillations shows up.  The dashed traces are the nominal error signals when I'm sitting at some CARM offset, and the solid traces are just after a step has been made. The glitch is only happening in the PRMI, not CARM and DARM. 

PRFPMI_currentErrSigs_25Nov2014.pdf

ELOG V3.1.3-